OpenStudy (anonymous):

Convert radian into degrees. Round your answers thousandths place. -0.48 I got -27.502 degrees.

OpenStudy (anonymous):

180/pie.x where x is angle in radian

OpenStudy (anonymous):

You put -0.48 x 180/pie which gives you -86.4/pie which gives you -27.502?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay ty

OpenStudy (unklerhaukus):

\[-0.48(\text{radiens})\]\[=-0.48(\text{radiens})\times1\]\[=-0.48(\text{radiens})\times\left(\frac{180°}{\pi(\text{radiens})}\right)\]\[= -0.48 \times 180°/\pi\]\[=-86.4°/\pi\]\[\approx-27.5°\]

OpenStudy (anonymous):

Thankyou!!!